### v1

@@ 5.1 Experimental Setting
Dataset: Gutenberg Project. The project Gutenberg [^6] https://www.gutenberg.org/ offers a high-quality open dataset of over 70,000 books, continuously expanding. PG-19 [23], a subset of 28,752 books extracted in 2019, has been included in RedPajama-Data [5] and The Pile [10] and used to train LLMs such as Pythia [1] and OpenLLaMA [11]). It is also widely used to evaluate MIAs (e.g. [19][18]). We use documents from Project Gutenberg in our experiments because of its quality, recognized relevance in MIA research and the availability of methods [19] to minimize bias.
@@
Bias assesment. The agnostic classifiers employed in this study utilize a Bayes algorithm for multinomially distributed data, focusing on the distribution of 1 to 3-grams. These classifiers are trained and applied using scikit-learn, chosen for its robustness in handling such data distributions. For the n-gram analysis, characters are treated as tokens when computing n-grams. We focus on n=7, as previous research [7] has shown that 7-grams reveal the most significant distributional differences. The n-gram analysis is conducted using a Bloom filter based on the implementation of [12], ensuring efficient and accurate bias detection.
@@
  • Min-kProb. This metric is based on the likelihood of the $k\%$ of tokens in a sequence $D$ that have the lowest probabilities, based on the preceding tokens [25].
  • Max-kProb. This is the inverse metric of Min-kProb, based on the tokens that have the highest probabilities. We use $k=10$ for both Min-kProb and Max-kProb.
  • zlib Ratio. This MIA identifies potential member when having a low ratio of the model’s perplexity to the entropy of the text [2]. This entropy is calculated as the number of bits required to compress the sequence using the zlib library [8].
  • Perplexity (ppl). This MIA leverages perplexity [2] as scores and then threshold them to classify samples as members or non-members.
  • Meta_MIA. This MIA is based on the work of [18], which aggregates 52 MIAs (including Min-kProb, perplexity, zlib Ratio, etc.) to create a single feature vector. A linear regressor is trained to learn the importance of weights for the different MIA attacks and thus classify their membership status.
@@ 5.2 Assuming No bias: Random Sample
By construction, $\mathcal{D}^{known}_{train}$ and $\mathcal{D}^{known}_{nm}$ are exempt of bias related to meta-data and time-shift. Since there is no reason to suspect a bias, we construct $\mathcal{D}^{NoB}_{train}$ and $\mathcal{D}^{NoB}_{nm}$ through a random sample. We compute the distribution of n-gram overlap of these two sets with reference to the left out members ($\mathcal{D}^{known}_{train}\backslash\mathcal{D}^{NoB}_{train}$) as described in Sec. 4.2. The resulting histograms are depicted in Fig. 1.
(a) Randomly selected members
(b) Randomly selected non-members
Figure 1: Histograms of n-gram overlap wrt left out members (KS-distance = 0.222)
@@
Conclusion.
The text of books written in the same time interval but added to project Gutenberg at different dates (before or after the extraction of PG-19) still exhibit n-gram shifts and a random sampling produce heavily biased datasets.
@@ 5.3 $No-Ngram$ to Minimize N-gram Bias
(a) $No-Ngram$ members
(b) $No-Ngram$ non-members
Figure 2: Histograms of n-gram overlap wrt left out members (KS-distance = 0.034)
We further assess residual bias by training a classifier on $\mathcal{D}^{NoB}_{train}$ and $\mathcal{D}^{NoB}_{nm}$. The ROC of each fold is illustrated in Fig. 3(a). As a reference, the evaluation of a classifier trained on randomly sampled datasets is depicted in Fig. 3(b). The agnostic classifier achieves on average over 5 folds 0.82 AUC ROC and 5%, 26%, and 49%TPR at 1%, 5%, and 10% FPR, respectively. This remains highly accurate and the accuracy loss is marginal when compared to random samples where an agnostic classifier achieves 0.84 AUC ROC and 3%, 30%, and 63% TPR at 1%, 5%, and 10% FPR, respectively.
(a) trained on $No-Ngram$ sets
(b) trained on randomly sampled set
Figure 3: ROC of 5-folds agnostic classifiers
Conclusion.
While the $No-Ngram$ algorithm has successfully reduced (if not eliminated) the bias in n-gram overlap, the produced sets can still be discriminated with high accuracy by an agnostic classifier. Contrarily to previous proposal [7], this suggests that distributional difference in n-gram overlap is insufficient as a metric of MIA benchmark difficulty.
@@ 5.4 Sampling unclassifiable datasets
The datasets being classifiable even when n-gram bias are minimized, we apply $No-Class$ [^7] Since we use a single classifier, we also ensure the same number of false positive, false negative, true positive and true negative in the selected sets. using a single classifier trained on randomly sampled set whose evaluation is presented in Fig. 3(b). To evaluate residual bias, we train a new agnostic classifier on the resulting sets whose evaluation is shown in Fig. 4.
Figure 4: ROC of 5-folds agnostic classifier trained on $No-Class$ sets
@@ 5.5 Assesment of MIAs
We assess 6 state of the art MIAs on 2 LLMs and the datasets random, $No-Ngram$, and $No-Class$ presented Sec. 5.25.3, and 5.4, respectively. Tables 1 and 2 report their TPR at 10%FPR and AUC ROC averaged over 5 runs.
@@ 6.0.1 Acknowledgements
This work was supported by the French grant iPoP PEPR (ANR-22-PECY-0002) and DATAIA. Jose Maria de Fuentes has also received support from the Spanish National Cybersecurity Institute (INCIBE) grant APAMciber within the framework of the Recovery, Transformation and Resilience Plan funds, financed by the European Union (Next Generation); and from UC3M’s Requalification programme, funded by the Spanish Ministerio de Ciencia, Innovacion y Universidades with EU recovery funds (Convocatoria de la Universidad Carlos III de Madrid de Ayudas para la recualificación del sistema universitario español para 2021-2023, de 1 de julio de 2021).

### v2

@@ 5.1 Experimental Setting
Dataset: Gutenberg Project. The project Gutenberg [^5] https://www.gutenberg.org/ offers a high-quality open dataset of over 70,000 books, continuously expanding. PG-19 [23], a subset of 28,752 books extracted in 2019, has been included in RedPajama-Data [5] and The Pile [10] and used to train LLMs such as Pythia [1] and OpenLLaMA [11]. It is also widely used to evaluate MIAs (e.g., [19][18]). We use documents from Project Gutenberg in our experiments because of its quality, recognized relevance in MIA research and the availability of methods [19] to minimize bias.
@@
Bias assessment. The agnostic classifiers employed in this study utilize a Bayes algorithm for multinomially distributed data, focusing on the distribution of 1 to 3-grams. These classifiers are trained and applied using scikit-learn, chosen for its robustness in handling such data distributions. For the n-gram analysis, characters are treated as tokens when computing n-grams. We focus on n=7, as previous research [7] has shown that 7-grams reveal the most significant distributional differences. The n-gram analysis is conducted using a Bloom filter based on the implementation of [12], ensuring efficient and accurate bias detection.
@@
  • Min-kProb is based on the likelihood of the $k\%$ of tokens in a sequence $D$ that have the lowest probabilities, based on the preceding tokens [25].
  • Max-kProb is the inverse metric of Min-kProb, based on the tokens that have the highest probabilities. We use $k=10$ for both Min-kProb and Max-kProb.
  • zlib Ratio identifies potential member when having a low ratio of the model’s perplexity to the entropy of the text [2]. This entropy is calculated as the number of bits required to compress the sequence using [8].
  • Perplexity (ppl) leverages perplexity [2] as scores and then threshold them to classify samples as members or non-members.
  • Meta_MIA is based on the work of [18], which aggregates 52 MIAs (including Min-kProb, perplexity, zlib Ratio, etc.) to create a single feature vector. A linear regressor is trained to learn the importance of weights for the different MIA attacks and thus classify their membership status.
@@ 5.2 Assuming No bias: Random Sample
By construction, $\mathcal{D}^{known}_{train}$ and $\mathcal{D}^{known}_{nm}$ are exempt of bias related to meta-data and time-shift. Since there is no reason to suspect a bias, we construct $\mathcal{D}^{Nob}_{train}$ and $\mathcal{D}^{Nob}_{nm}$ through a random sample. We compute the distributions of n-gram overlap of these two sets with reference to the left out members ($\mathcal{D}^{known}_{train}\backslash\mathcal{D}^{Nob}_{train}$) as described in Sec. 4.2. The result is depicted as histograms in Fig. 1.
(a) Randomly selected members.
(b) Randomly selected non-members.
Figure 1: Histograms of n-gram overlap w.r.t. left out members (KS-dist = 0.222).
@@
Conclusion. The text of books written in the same time interval but added to project Gutenberg at different dates (before or after the extraction of PG-19) still exhibit n-gram shifts and a random sampling produce heavily biased datasets.
@@ 5.3 $No-Ngram$ to Minimize N-gram Bias
(a) $No-Ngram$ members.
(b) $No-Ngram$ non-members.
Figure 2: Histograms of n-gram overlap w.r.t. left out members (KS-dist = 0.034).
(c) Trained on $No-Ngram$ sets.
(d) Trained on randomly sampled set.
Figure 3: ROC of 5-folds agnostic classifiers.
We further assess residual bias by training a classifier on $\mathcal{D}^{Nob}_{train}$ and $\mathcal{D}^{Nob}_{nm}$. The ROC of each fold is illustrated in Fig. 2(c). As a reference, the evaluation of a classifier trained on randomly sampled datasets is depicted in Fig. 2(d). The agnostic classifier achieves on average over 5 folds 0.82 AUC ROC and 5%, 26%, and 49%TPR at 1%, 5%, and 10% FPR, respectively. This remains highly accurate and the accuracy loss is marginal when compared to random samples where an agnostic classifier achieves 0.84 AUC ROC and 3%, 30%, and 63% TPR at 1%, 5%, and 10% FPR, respectively.
Conclusion. While the $No-Ngram$ algorithm has successfully reduced (if not eliminated) the bias in n-gram overlap, the produced sets can still be discriminated with high accuracy by an agnostic classifier. Contrarily to previous proposal [7], this suggests that distributional difference in n-gram overlap is insufficient as a metric of MIA benchmark difficulty.
@@ 5.4 Sampling Unclassifiable Datasets
The datasets being classifiable even when n-gram bias are minimized, we apply $No-Class$ [^6] Since we use a single classifier, we also ensure the same number of false positive, false negative, true positive and true negative in the selected sets. using a single classifier trained on randomly sampled set whose evaluation is presented in Fig. 2(d). To evaluate residual bias, we train a new agnostic classifier on the resulting sets whose evaluation is shown in Fig. 4.
Figure 4: ROC of 5-folds agnostic classifier trained on $No-Class$ sets.
@@ 5.5 Assesment of MIAs
We assess 6 state of the art MIAs on 2 LLMs and the datasets random, $No-Ngram$, and $No-Class$ presented in Sec. 5.25.3, and 5.4, respectively. Tables 2 and 2 report their TPR at 10%FPR and AUC ROC averaged over 5 runs.
@@ 6.0.1 Acknowledgements
This work was supported by the “ANR 22-PECY-0002” IPoP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR and DATAIA. Jose Maria de Fuentes has also received support from the Spanish National Cybersecurity Institute (INCIBE) grant APAMciber within the framework of the Recovery, Transformation and Resilience Plan funds, financed by the European Union (Next Generation); and from UC3M’s Requalification programme, funded by the Spanish Ministerio de Ciencia, Innovacion y Universidades with EU recovery funds (Convocatoria de la Universidad Carlos III de Madrid de Ayudas para la recualificación del sistema universitario español para 2021-2023, de 1 de julio de 2021).